# OMPGPT: A Generative Pre-trained Transformer Model for OpenMP

### v2

@@ Abstract
Large language models (LLMs)such as ChatGPT have significantly advanced the field of Natural Language Processing (NLP). This trend led to the development of code-based large language models such as StarCoder, WizardCoder, and CodeLlama, which are trained extensively on vast repositories of code and programming languages. While the generic abilities of these code LLMs are useful for many programmers in tasks like code generation, the area of high-performance computing (HPC) has a narrower set of requirements that make a smaller and more domain-specific model a smarter choice. This paper presents OMPGPT, a novel domain-specific model meticulously designed to harness the inherent strengths of language models for OpenMP pragma generation. Furthermore, we leverage prompt engineering techniques from the NLP domain to create Chain-of-OMP, an innovative strategy designed to enhance OMPGPT’s effectiveness. Our extensive evaluations demonstrate that OMPGPT outperforms existing large language models specialized in OpenMP tasks and maintains a notably smaller size, aligning it more closely with the typical hardware constraints of HPC environments. We consider our contribution as a pivotal bridge, connecting the advantage of language models with the specific demands of HPC tasks.
@@ 2.1 Generative Pre-trained Transformers and Code LLMs
StarCoder [17] is a 15B parameter model trained for code generation or completion. The training dataset, the Stack [15], has 1 trillion tokens sourced from a large collection of permissively licensed GitHub repositories. CodeLlama [22] is a code-generating LLM based on Llama2 [24] by Meta, specialized for code by training with code-specific datasets. It has parameter sizes of 7B, 13B, and 34B, trained with 500B tokens of code data. Most code LLMs (e.g., StarCoder) have been trained on raw code data without instruction fine-tuning. However, the recent WizardCoder [20], enables Code LLMs with instruction fine-tuning by adapting the Evol-Instruct methods for coding tasks, and it was shown to improve the performance of code generation [4].
@@ 3.2 OMPGPT Training & Inference
Dataset. HPCorpus, introduced by Kadosh et al. [13], is an extensive HPC database derived from GitHub repositories containing C, C++, and Fortran code. Notably, approximately 45% of the repositories use some form of parallel programming. The primary parallel programming mode is OpenMP, making HPCorpus suitable for training models for OpenMP tasks. We use C and C++ code from HPCorpus for our model’s training and fine-tuning, with 144,522 C code repositories and 150,481 C++ code repositories, totaling 8,781,759 C code functions and 68,233,984 C++ code functions (total 72.39 GB). Figure LABEL:fig:data_dist shows the distribution of OpenMP pragmas in the HPCorpus OpenMP subset used for OMPGPT fine-tuning. We allocate 10% of the data as a test set for both training and fine-tuning to maintain evaluation integrity.
@@
    Filtering code by size: Following practices used in previous research such as PolyCoder [28], we filter out large code segments (more than 100 tokens or greater than 1 MB) from HPCorpus.
3.2.1 Compute Hardware.
In all experiments, we utilized a single node on an HPC cluster, equipped with a dual-socket AMD EPYC Rome 7402 (24 cores/socket) and 512 GB DDR4-3200 RAM, along with 4 x NVIDIA A100 40GB HBM2e Tensor Core GPUs connected via NVLink3 to each other.
@@ 4.3 OpenMP Pragma Generation using Chain-of-OMP
Table 2: Chain-of-OMP evaluation with OMPGPT.
@@ 5 Related Work
Chain-of-thought prompting improves LM reasoning ability, showing promise on math word problems, commonsense reasoning [27], and summarizing software components [23]. This prompting strategy has also enhanced code generation from models like ChatGPT [16, 18] and bug reproduction from reports [12]. Most relevantly, it has been applied to data race detection [6] and compilable code generation [5].

### v3

@@ Abstract
Large language models (LLMs)such as ChatGPT have significantly advanced the field of Natural Language Processing (NLP). This trend led to the development of code-based large language models such as StarCoder, WizardCoder, and CodeLlama, which are trained extensively on vast repositories of code and programming languages. While the generic abilities of these code LLMs are helpful for many programmers in tasks like code generation, the area of high-performance computing (HPC) has a narrower set of requirements that make a smaller and more domain-specific model a smarter choice. This paper presents OMPGPT, a novel domain-specific model meticulously designed to harness the inherent strengths of language models for OpenMP pragma generation. Furthermore, we leverage prompt engineering techniques from the NLP domain to create Chain-of-OMP, an innovative strategy designed to enhance OMPGPT’s effectiveness. Our extensive evaluations demonstrate that OMPGPT outperforms existing large language models specialized in OpenMP tasks and maintains a notably smaller size, aligning it more closely with the typical hardware constraints of HPC environments. We consider our contribution as a pivotal bridge, connecting the advantage of language models with the specific demands of HPC tasks.
@@ 2.1 Generative Pre-trained Transformers and Code LLMs
StarCoder [15] is a 15B parameter model trained for code generation or completion. The training dataset, the Stack [13], has 1 trillion tokens sourced from a large collection of permissively licensed GitHub repositories. CodeLlama [20] is a code-generating LLM based on Llama2 [22] by Meta, specialized for code by training with code-specific datasets. It has parameter sizes of 7B, 13B, and 34B, trained with 500B tokens of code data. Most code LLMs (e.g., StarCoder) have been trained on raw code data without instruction fine-tuning. However, the recent WizardCoder [18], enables Code LLMs with instruction fine-tuning by adapting the Evol-Instruct methods for coding tasks, and it was shown to improve the performance of code generation.
@@ 3.2 OMPGPT Training & Inference
Dataset. HPCorpus, introduced by Kadosh et al. [10], is an extensive HPC database derived from GitHub repositories containing C, C++, and Fortran code. Notably, approximately 45% of the repositories use some form of parallel programming. The primary parallel programming mode is OpenMP, making HPCorpus suitable for training models for OpenMP tasks. We use C and C++ code from HPCorpus for our model’s training and fine-tuning, with 144,522 C code repositories and 150,481 C++ code repositories, totaling 8,781,759 C code functions and 68,233,984 C++ code functions (a total of 72.39 GB). Figure 3 shows the distribution of OpenMP pragmas in the HPCorpus OpenMP subset used for OMPGPT fine-tuning. We allocate 10% of the data as a test set for both training and fine-tuning to maintain evaluation integrity.
@@
    Filtering code by size: Following practices used in previous research, such as PolyCoder [26], we filter out large code segments (more than 100 tokens or greater than 1 MB) from HPCorpus.
3.2.1 Hardware.
In all experiments, we utilized a single node on an HPC cluster equipped with a dual-socket AMD EPYC Rome 7402 (24 cores/socket) and 512 GB DDR4-3200 RAM, along with 4 x NVIDIA A100 40GB HBM2e Tensor Core GPUs connected via NVLink3 to each other.
@@ 4.3 OpenMP Pragma Generation using Chain-of-OMP
Table 2: Chain-of-OMP effectiveness evaluation with OMPGPT.
@@ 5 Related Work
Chain-of-thought prompting improves LM reasoning ability, showing promise on math word problems, commonsense reasoning [25], and summarizing software components [21]. This prompting strategy has also enhanced code generation from models like ChatGPT [14, 16] and bug reproduction from reports [9]. Most relevantly, it has been applied to correcting compilable code [3].
@@ 6 Conclusion and Future Work
7 Acknowledgement
This project was funded by NSF (#$2211982$) and Intel Labs. We would like to thank them for their generous support. Additionally, we extend our gratitude to the Research IT team [^1] https://researchit.las.iastate.edu/ of Iowa State University for their continuous support in providing access to HPC clusters for conducting the experiments of this project.